OpenStudy (anonymous):

how do i factor 2y+30=2.....the answer they are looking for is 2y+30=2(?)

OpenStudy (anonymous):

2(y+15) = 2y + 30 ? = y + 15

OpenStudy (anonymous):

all it says is complete factoring...and wants me to fill the blnk in of 2y+30=2( )..supposed to fill the inside of the parenthesis...